# thanks: Deceasedthanks: Deceased

### v1

@@ Abstract
Hypothetical axions provide a compelling explanation for dark matter and could be emitted from the hot solar interior. The CERN Axion Solar Telescope (CAST) has been searching for solar axions via their back conversion to X-ray photons in a 9-T 10-m long magnet directed towards the Sun. We report on an extended run with the IAXO (International Axion Observatory) pathfinder detector, doubling the previous exposure time. The detector was operated with a xenon-based gas mixture for part of the new run, providing technical insights for future detector configurations in IAXO. No counts are detected in the 95% signal-encircling region during the new run, while one is expected. The new data improve the axion-photon coupling limit to 5.7×$\times 10^{-11}\,$eV-1 at 95% C.L., the most restrictive experimental limit to date.
Introduction.—Very light pseudoscalar bosons, generically called axion-like particles (ALPs), appear in many motivated extensions of the Standard Model [Ringwald:2012hr, Jaeckel:2010ni]. The paradigmatic example in this category is the axion, whose existence follows from the Peccei–Quinn mechanism as an explanation of why QCD (quantum chromodynamics) is perfectly time-reversal invariant within current experimental precision [Peccei:1977hh, Wilczek:1977pj, Weinberg:1977ma]. Axions and ALPs can be dark matter in the form of classical field oscillations that were excited in the early universe, by the re-alignment mechanism [Preskill:1982cy, Abbott:1982af, Dine:1982ah], or by the decay of topological defects of the axion field [Marsh:2015xka].
@@
where $a$ is the axion field, $F$ the electromagnetic field-strength tensor, $g_{a\gamma}$ the coupling constant, ${\bf E}$ the electric field and ${\bf B}$ the magnetic field. This vertex enables the decay $a\to\gamma\gamma$, as well as the Primakoff production in stars, i.e., the $\gamma\to a$ scattering in the Coulomb fields of charged particles in the stellar plasma, and the coherent conversion $a\leftrightarrow\gamma$ in laboratory or astrophysical $B$-fields [Sikivie:1983ip, Raffelt:1987im].
@@
Axion helioscopes make use of a dipole magnet directed at the Sun to convert axions to X-rays. Solar axions can be produced in several processes, depending on their model-dependent interaction channels. We specifically consider axion production by Primakoff scattering of thermal photons deep in the Sun, a process that depends on the coupling constant $g_{a\gamma}$, which is also used for detection.
@@
This detection concept has been followed by the CERN Axion Solar Telescope (CAST), the most powerful axion helioscope so far built [Anastassopoulos:2017ftl]. CAST was in operation at CERN from 2003 until 2021. During this time, the experiment has gone through different phases and has released a number of results, including a first phase using evacuated magnet bores [Zioutas:2004hi, Andriamonje:2007ew], followed by “gas phases” with 4He [Arik:2008mq, Arik:2015cjv] and 3He [Aune:2011rx, Arik:2013nya], to cover sensitivity to higher $m_{a}$ values. Later on, CAST returned to evacuated magnet bores but with an improved detection line, dubbed IAXO (International Axion Observatory) pathfinder, that combined a new Micromegas detector with lower background levels, as well as a new X-ray telescope built specifically for axion searches [Aznar:2015iia] that is based on technology developed for the Nuclear Spectroscopic Telescope Array (NuSTAR) [nustar2013]. This allowed CAST to produce what is at present the strongest experimental upper bound to the solar axion-photon coupling [Anastassopoulos:2017ftl].
@@
CAST also produced constraints to other (non-Primakoff) axion/ALP production channels in the Sun, [Andriamonje:1186141, Barth:2013sma, CAST:2009klq], as well as to chameleons [Anastassopoulos2015] and hidden photons [Redondo:2015iea]. In a later stage, CAST expanded its scope to the search for dark matter axions [CAST:2020rlf, CAST-CAPP:2022] and solar chameleons via pressure sensing [CAST-KWISP:2019].
@@
In this last phase, the IAXO pathfinder line kept operating from September 2019 to June 2021, in order to improve statistics of the 2017 result. As a novelty, part of this data was taken with a new gas recirculation and filtering system to use a Xe-based gas mixture for the Micromegas detectors. The goal was to get lower background for the science analysis by testing this conversion gas in real experiment conditions, which will be valuable for future Xe-based detectors. The analysis of the data taken in these runs is the main result of the present article. These new results are combined with the latest GridPix data from October 2017 to December 2018. The GridPix detector [Krieger:2013cfa] is optimized for lower energy signals (i.e. it is more sensitive to axion-electron coupling $g_{ae}$) due to a lower energy threshold than Micromegas detectors, and its results will be detailed in a forthcoming publication currently in preparation. Here these data only improved the result marginally in the upper limit for $g_{a\gamma}$.
@@
In CAST, argon mixtures were used historically as active gas in the chamber due to being well studied and posing little engineering challenges [Barth:2013sma, Anastassopoulos:2017ftl, SebastianTesis], and it was also used during the first period of the data taking discussed in this work. However, argon’s X-ray fluorescence around 3 keV results in an increased background level at the energy range where the maximum of solar axion Primakoff flux is expected. Consequently, there has been a shift towards exploring alternative gas mixtures, particularly those based on xenon. Unlike the argon mixtures, which were utilized in an open loop system, xenon mixtures were employed in a closed loop system. Special attention has been given to potential contamination of water vapor or oxygen from leaks or outgassing. Implementing this change required the development of a new and more sophisticated gas recirculation system, incorporating moisture and oxygen filters, a recirculation pump and a buffer volume. The gas mixtures used during the last data taking campaign were Ar + 2.3% isobutane at 1.4 bar and 48.85%Xe + 48.85%Ne + 2.3% isobutane at 1.05 bar. Xenon-based gas mixtures have a higher detection efficiency, allowing for the use of lower pressures. We demonstrate that their use is possible, paving the way for the future use of even thinner windows and reduced gas pressures.
Data taking.—Calibration of the detector was performed at different energies in the CAST X-ray tube at CERN using both Ar- and Xe-based mixtures, providing information on its response to photons of different energies that was later used for event discrimination. During the data taking periods at CAST, calibrations with a 55Fe source were taken on a daily basis. This enabled the calibration of each background and tracking run in energy and the evaluation of the detector performance, stability and energy threshold over time. Once the detector was placed in CAST, the correct alignment of the full beamline was verified by placing a piezoelectric X-ray source on the opposite side of the magnet. In addition to these measurements, targets were securely positioned during geometric surveys, allowing for verification of the detector’s position. Furthermore, the Sun can be filmed twice per year with an optical telescope and a camera attached to the magnet, ensuring the pointing accuracy of the setup.
@@
Table 1: Datasets of the presented data taking campaign. Ar: Ar + 2.3% isobutane at 1.4 bar; Xe: 48.85% Xe + 48.85% Ne + 2.3% isobutane at 1.05 bar. The software efficiency $\epsilon_{s}$ at 5.9 keV is indicated for each dataset.
@@
The detector was stable during most of the data taking campaign and showed an energy resolution of 20 to 23% at 5.9  keV. For datasets 1 and 2 the gain variation was 3.3%, with an energy threshold consistently below 0.4 keV. Dataset 3 required more attention to maintain optimal gas gain and transparency due to the recirculation system and its effect on the gas quality. Consequently, voltage and flow parameters were dynamically adjusted as required. The range of gain variation was 2.5% to 9%. Since the gain changes are constantly monitored they can be easily corrected with every calibration and they do not affect the data quality and analysis. The energy threshold remained below 0.75 keV for most (83%) of the runs. A more detailed overview of the data taking stability is provided in the Supplemental Material.
Data analysis and results.— The data analysis and background rejection was performed using the REST-for-physics framework [REST, REST-git], a ROOT-based collaborative software developed for data analysis and Geant4-based [AGOSTINELLI2003250] simulations of rare event searches experiments and gaseous detectors using TPCs. This analysis was performed in a fully blind manner to prevent bias in the results.
The data analysis chain turns the raw data into events with a given energy and a physical position on the readout plane, for which topological observables can be computed. This allows for powerful background rejection as X-rays (small, symmetric, point-like and single-track events) are easily identified. Furthermore, events that occur in coincidence with a cosmic event in the active shielding are also removed. See [MMcuts] and also the Supplemental Material for an overview of the Micromegas events analysis and definition of the X-ray selection algorithms.
@@
This background level is compared to the rate of X-ray-like observed events during the 314.6 h of solar tracking time. The same selection algorithms are applied to these data and the aim is to look for any X-ray event excess during axion sensitive conditions. In order to do so, ray tracing simulations provide the information of the signal spot shape and position on the readout plane. The signal probability density function (PDF) is energy dependent, but most of the flux is in all cases focused into an area of a few mm2. The ray-tracing simulations were performed for a solar axion flux and also for an X-ray source such as the one used for alignment calibrations. Comparing the latter with the real data allows to align the simulations density contours with the experimental data (Fig. 2). For this purpose, the contours containing 68%, 85%, 95% and 99% of the signal were computed. The same alignment was applied to the X-ray events during tracking (Fig. 3), where we observe 0 counts in the energy range of interest (2 to 7 keV) in the 95% signal encircling region in all 3 datasets, while expectations based on the background rate and exposure time is 1 count.
@@
Limit on the axion-photon coupling.— We use an unbinned likelihood method to compute an upper limit on $g_{a\gamma}$, following the same methodology as in [Aune:2011rx, Arik:2013nya, Arik:2015cjv, Anastassopoulos:2017ftl], as it is better suited for low count experiments. The defined likelihood function is
@@
where the detector response as a function of energy is encoded in the $\epsilon(E,\vec{x_{i}})$ term, which includes the X-ray optics efficiency of the telescope, the hardware and software efficiencies, and the axion signal simulations defining the expected signal distribution (Fig. 3). The axion to photon conversion probability in an homogeneous magnetic field $B$ of length $L$ is
@@
where $q=m_{a}^{2}/2E$ is the momentum transfer between axion and photon in vacuum. Finally, the differential Primakoff solar axion flux in keV-1 cm-2 s-1 is given by the expression [Andriamonje:2007ew]
@@
where $g_{10}=g_{a\gamma}/(10^{-10}\,\mathrm{GeV}^{-1}).$
@@
The resulting PDF is combined with results from [Anastassopoulos:2017ftl], which uses data up to 2015 and serves as the benchmark for the axion-photon coupling set by CAST. Additionally, data from the 2017-2018 campaign with a GridPix detector [Krieger:2013cfa] [SebastianTesis], providing an extra 160 hours of data, is included. This combination allowed us to create an overall PDF that encapsulates all CAST data with sensitivity to $g_{a\gamma}$.
@@
For higher axion masses up to 0.06 keV we have also set the most stringent limit, getting closer to the QCD axion band. For $m_{a}\gtrsim 0.06$ eV, the data taken during the 4He and 3He as buffer gas periods is still the most competitive result from CAST [Zioutas:2004hi, Andriamonje:2007ew, Arik:2008mq]. The uncertainty in this latest result is dominated by the statistical effect of the low count rate available. The effect of systematic errors on this limit is discussed in the Supplemental Material, but remains well below 10% of the statistical uncertainty of the result.
Figure 4: Parameter space for axions and ALPs showing the latest constraints in $g_{a\gamma}$. The red line indicates the region excluded by this work, which builds upon the former limit indicated in blue. The yellow band represents the region QCD axion models point to.
This result is a mild statistical underfluctuation as compared to the expected sensitivity of the experiment, defined as the median upper limit of multiple Monte Carlo simulations, which for the current exposure time is $g_{a\gamma}<0.59\times 10^{-10}$ GeV-1. See the Supplemental Material for a more detailed explanation of the simulations of the experiment’s potential sensitivity.
Conclusions.— The result presented in this paper represents the new best limit on $g_{a\gamma}$ in the range $m_{a}\lesssim 0.02$ eV, superseding our previous best result, and going beyond the limit derived from the energy loss of horizontal branch (HB) stars  [Ayala:2014pea, Straniero:2015nvc]. Currently, our bound is surpassed only by astrophysical considerations on the axion impact on the R2 parameter in globular clusters, which lead to $g_{a\gamma}<0.47\times 10^{-10}\mathrm{GeV}^{-1}$. This last result, however, relies on the accurate counting of stars in globular clusters as well as on numerical simulations of stars in late evolutionary stages, and is marred by considerably larger uncertainties. This improvement originates not only from the additional statistics, but also from an improved detection efficiency, especially at low energies. This is in part thanks to the use of a Xe-based gas mixture in the conversion volume, something that also improves the background in the region of interest. This aspect constitutes a relevant technical achievement that provides useful operational experience for future similar implementations in BabyIAXO and IAXO.
@@
Acknowledgments.— We thank CERN for hosting the experiment and for technical support to operate the magnet and cryogenics. We thank V. Burwitz, G. Hartner of the MPE PANTER X-ray test facility for providing the opportunity to calibrate the X-ray telescope and for assistance in collecting and analysing the characterization data. We acknowledge support from NSERC (Canada); MSE (Croatia); CEA (France); BMBF (Germany) under the grant numbers 05 CC2EEA/9 and 05 CC1RD1/0, DFG (Germany) under grant numbers HO 1400/7-1 and EXC-153, GSRT (Greece), NSRF: Heracleitus II; Agence Nationale de la Recherche (France) ANR-19-CE31-133 0024; the Spanish Agencia Nacional de Investigación (AEI) under grants PID2019-108122GB-C31 and PID2022-137268NB-C51,funded by MCIN/AEI/10.13039/501100011033/FEDER, as well as the “Planes complementarios, Programa de Astrofísica y Física de Altas Energías” from the “European Union NextGenerationEU/PRTR” funds; the European Research Council (ERC) under grants ERC-2017-AdG 788781 (IAXO+) and ERC 802836; Turkish Atomic Energy Authority (TAEK); NSF (USA) under Award No. 0239812; NASA under the grant number NAG5-10842; and IBS (Korea) with code IBS-R017-D1. Part of this work was performed under the auspices of the US Department of Energy by Lawrence Livermore National Laboratory under Contract No. DE-AC52-07NA27344 and funded through Lawrence Livermore National Laboratory Directed Research and Development project LDRD-17-ERD-030.

### v3

@@ Abstract
Hypothetical axions provide a compelling explanation for dark matter and could be emitted from the hot solar interior. The CERN Axion Solar Telescope (CAST) has been searching for solar axions via their back conversion to X-ray photons in a 9-T 10-m long magnet directed towards the Sun. We report on an extended run with the IAXO (International Axion Observatory) pathfinder detector, doubling the previous exposure time. The detector was operated with a xenon-based gas mixture for part of the new run, providing technical insights for future detector configurations in IAXO. No counts are detected in the 95% signal-encircling region during the new run, while 0.75 are expected. The new data improve the axion-photon coupling limit to 5.8×$\times 10^{-11}\,$eV-1 at 95% C.L. (for $m_{a}\lesssim 0.02$ eV), the most restrictive experimental limit to date.
Introduction.—Very light pseudoscalar bosons, generically called axion-like particles (ALPs), appear in many motivated extensions of the Standard Model [1, 2]. The paradigmatic example in this category is the axion, whose existence follows from the Peccei–Quinn mechanism as an explanation of why QCD (quantum chromodynamics) is perfectly time-reversal invariant within current experimental precision [3, 4, 5]. Axions and ALPs can be dark matter in the form of classical field oscillations that were excited in the early universe, by the re-alignment mechanism [6, 7, 8], or by the decay of topological defects of the axion field [9].
@@
where $a$ is the axion field, $F$ the electromagnetic field-strength tensor, $g_{a\gamma}$ the coupling constant, ${\bf E}$ the electric field and ${\bf B}$ the magnetic field. This vertex enables the decay $a\to\gamma\gamma$, as well as the Primakoff production in stars, i.e., the $\gamma\to a$ scattering in the Coulomb fields of charged particles in the stellar plasma, and the coherent conversion $a\leftrightarrow\gamma$ in laboratory or astrophysical $B$-fields [11, 16].
@@
Solar axions can be produced in several processes, depending on their model-dependent interaction channels. We specifically consider axion production by Primakoff scattering of thermal photons deep in the Sun, a process that depends on the coupling constant $g_{a\gamma}$, which is also used for detection. Following this principle, axion helioscopes make use of a dipole magnet directed at the Sun to convert axions to X-rays.
@@
This detection concept has been followed by the CERN Axion Solar Telescope (CAST), the most powerful axion helioscope so far built [17]. CAST was in operation at CERN from 2003 until 2021. During this time, the experiment has gone through different phases and has released a number of results, including a first phase using evacuated magnet bores [18, 19], followed by “gas phases” with 4He [20, 21] and 3He [22, 23], to cover sensitivity to higher $m_{a}$ values. Later on, CAST returned to evacuated magnet bores but with an improved detection line, dubbed IAXO (International Axion Observatory) pathfinder, that combined a new Micromegas detector with lower background levels, as well as a new X-ray telescope built specifically for axion searches [24] that is based on technology developed for the Nuclear Spectroscopic Telescope Array (NuSTAR) [25]. This allowed CAST to produce what is at present the strongest experimental upper bound to the solar axion-photon coupling [17].
@@
CAST also produced constraints to other (non-Primakoff) axion/ALP production channels in the Sun, [26, 27, 28], as well as to chameleons [29] and hidden photons [30]. In a later stage, CAST expanded its scope to the search for dark matter axions [31, 32] and solar chameleons via pressure sensing [33].
@@
In this last phase, the IAXO pathfinder line kept operating from September 2019 to June 2021, in order to improve statistics of the 2017 result. As a novelty, part of this data was taken with a new gas recirculation and filtering system to use a Xe-based gas mixture for the Micromegas detectors. The goal was to get lower background for the science analysis by testing this gas filling the Micromegas detector volume in real experiment conditions, which will be valuable for future Xe-based detectors. The analysis of the data taken in these runs is the main result of the present article. These new results are combined with the latest GridPix data from October 2017 to December 2018. The GridPix detector [34] is optimized for lower energy signals (i.e. it is more sensitive to axion-electron coupling $g_{ae}$) due to a lower energy threshold than Micromegas detectors as GridPix can detect individual electrons, and its results will be detailed in a forthcoming publication currently in preparation. Here these data only improved the result marginally in the upper limit for $g_{a\gamma}$.
@@
In CAST, argon mixtures were used historically as gas in the Micromegas chamber due to being well studied and posing little engineering challenges [27, 17, 39], and it was also used during the first period of the data taking discussed in this work. However, argon’s X-ray fluorescence around 3 keV results in an increased background level at the energy range where the maximum of solar axion Primakoff flux is expected. Consequently, there has been a shift towards exploring alternative gas mixtures, particularly those based on xenon. Unlike the argon mixtures, which were utilized in an open loop system, xenon mixtures were employed in a closed loop system. Special attention has been given to potential contamination of water vapor or oxygen from leaks or outgassing. Implementing this change required the development of a new and more sophisticated gas recirculation system, incorporating moisture and oxygen filters, a recirculation pump and a buffer volume. The gas mixtures used during the last data taking campaign were Ar + 2.3% isobutane at 1.4 bar and 48.85%Xe + 48.85%Ne + 2.3% isobutane at 1.05 bar. Xenon-based gas mixtures have a higher detection efficiency, allowing for the use of lower pressures that reduce the pressure difference between the vacuum pipe and the Micromegas gas volume, which would allow the use of thinner windows with a higher X-ray transparency, increasing the overall efficiency. We demonstrate that their use is possible, paving the way for the future use of even thinner windows and reduced gas pressures.
Data taking.—Calibration of the detector was performed at different energies in the CAST X-ray tube at CERN using both Ar- and Xe-based mixtures, providing information on its response to photons of different energies that was later used for event discrimination. During the data taking periods at CAST, calibrations using the 5.9 keV peak of a 55Fe source were taken on a daily basis. This enabled the calibration of each background and tracking run in energy and the evaluation of the detector performance, stability and energy threshold over time. Once the detector was placed in CAST, the correct alignment of the full beamline was verified by placing a $\sim 3$ keV X-ray generator on the opposite side of the magnet. In addition to these measurements, targets were securely positioned during geometric surveys, allowing for verification of the detector’s position. Furthermore, the Sun can be filmed twice per year with an optical telescope and a camera attached to the magnet, ensuring the pointing accuracy of the setup.
@@
Table 1: Datasets of the presented data taking campaign. Ar: 97.7% Ar + 2.3% isobutane at 1.4 bar; Xe: 48.85% Xe + 48.85% Ne + 2.3% isobutane at 1.05 bar. The software efficiency $\epsilon_{s}$ at 5.9 keV is indicated for each dataset.
@@
The detector was stable during most of the data taking campaign and showed an energy resolution of 20 to 23% at 5.9  keV. For datasets 1 and 2 the gain variation was 3.3%, with an energy threshold consistently below 0.4 keV. Dataset 3 required more attention to maintain optimal gas gain and transparency due to the recirculation system and its effect on the gas quality. Consequently, voltage and flow parameters were dynamically adjusted as required. The range of gain variation was 2.5% to 9%. Since the gain changes are constantly monitored they can be easily corrected with every calibration and they do not affect the data quality and analysis. The energy threshold remained below 0.75 keV for most (83%) of the runs. A more detailed overview of the data taking stability is provided in the supplemental material [36].
Data analysis and results.— The data analysis and background rejection was performed using the REST-for-physics framework [40, 41], a ROOT-based collaborative software developed for data analysis and Geant4-based [42] simulations of rare event searches experiments and gaseous detectors using TPCs. This analysis was performed in a fully blind manner to prevent bias in the results.
The data analysis chain turns the raw data into events with a given energy and a physical position on the readout plane, for which topological observables can be computed. This allows for powerful background rejection as X-rays (small, symmetric, point-like and single-track events) are easily identified. Furthermore, events that occur in coincidence with a cosmic event in the active shielding are also removed. See [43] and also the Supplemental Material [36] for an overview of the Micromegas events analysis and definition of the X-ray selection algorithms.
@@
This background level is compared to the rate of X-ray-like observed events during the 314.6 h of solar tracking time. The same selection algorithms are applied to these data and the aim is to look for any X-ray event excess during axion sensitive conditions. In order to do so, ray tracing simulations provide the information of the signal spot shape and position on the readout plane. The signal probability density function (PDF) is energy dependent, but most of the flux is in all cases focused into an area of a few mm2. The ray-tracing simulations were performed for a solar axion flux originating from an extended source with and angular size equivalent to the solar core, placed at infinity. Simulations were also conducted for an X-ray source like the one used for alignment calibrations, originating from a point source placed at  10 m distance. Comparing the latter with the real data allows to align the simulations density contours with the experimental data (Fig. 2). For this purpose, the contours containing 68%, 85%, 95% and 99% of the signal were computed. The same alignment was applied to the X-ray events during tracking (Fig. 3), where we observe 0 counts in the energy range of interest (2 to 7 keV) in the 95% signal encircling region in all 3 datasets, while expectations based on the background rate and exposure time is 0.75 counts.
@@
Limit on the axion-photon coupling.— We use an unbinned likelihood method to compute an upper limit on $g_{a\gamma}$, following the same methodology as in [22, 23, 21, 17], as it is better suited for low count experiments. The defined likelihood function is
@@
where the detector response as a function of energy is encoded in the $\epsilon(E,\vec{x_{i}})$ term, which includes the X-ray optics efficiency of the telescope, the hardware and software efficiencies, and the axion signal simulations defining the expected signal distribution (Fig. 3). The axion to photon conversion probability in an homogeneous magnetic field $B$ of length $L$ in vacuum is
@@
where $q=m_{a}^{2}/2E$ is the momentum transfer between axion and photon in vacuum. Finally, the differential Primakoff solar axion flux in keV-1 cm-2 s-1 is given by the expression [19]
@@
where $g_{10}=g_{a\gamma}/(10^{-10}\,\mathrm{GeV}^{-1})$ and energy $E$ is in keV.
@@
The resulting PDF is combined with results from [17], which uses data up to 2015 and serves as the benchmark for the axion-photon coupling set by CAST. Additionally, data from the 2017-2018 campaign with a GridPix detector [34] [39], providing an extra 160 hours of data, is included. This combination allowed us to create an overall PDF that encapsulates all CAST data with sensitivity to $g_{a\gamma}$.
@@
For higher axion masses up to 0.06 keV we have also set the most stringent limit, getting closer to the QCD axion band. For $m_{a}\gtrsim 0.06$ eV, the data taken during the 4He and 3He as buffer gas periods is still the most competitive result from CAST [18, 19, 20]. The uncertainty in this latest result is dominated by the statistical effect of the low count rate available. The effect of systematic errors on this limit is discussed in the Supplemental Material [36], which includes references [44, 45], but remains well below 10% of the statistical uncertainty of the result.
Figure 4: Parameter space for axions and ALPs showing the latest constraints in $g_{a\gamma}$. The red line indicates the region excluded by this work, which builds upon the former limit indicated in black. The yellow band represents the region QCD axion models point to.
This result is a mild statistical underfluctuation as compared to the expected sensitivity of the experiment, defined as the median upper limit of multiple Monte Carlo simulations, which for the current exposure time is $g_{a\gamma}<0.59\times 10^{-10}$ GeV-1. See the Supplemental Material [36], which includes reference [46], for a more detailed explanation of the simulations of the experiment’s potential sensitivity.
Conclusions.— The result presented in this paper represents the new best limit on $g_{a\gamma}$ in the range $m_{a}\lesssim 0.02$ eV, superseding our previous best result, and going beyond the limit derived from the energy loss of horizontal branch (HB) stars  [47, 48]. Currently, our bound is surpassed only by astrophysical considerations on the axion impact on the R2 parameter in globular clusters, which lead to $g_{a\gamma}<0.47\times 10^{-10}\mathrm{GeV}^{-1}$. This last result, however, relies on the accurate counting of stars in globular clusters as well as on numerical simulations of stars in late evolutionary stages, and is marred by considerably larger uncertainties. This improvement originates not only from the additional statistics, but also from an improved detection efficiency, especially at low energies. This is in part thanks to the use of a Xe-based gas mixture in the Micromegas detector gas volume, something that also improves the background in the region of interest. This aspect constitutes a relevant technical achievement that provides useful operational experience for future similar implementations in BabyIAXO and IAXO.
@@
Acknowledgments.— We thank CERN for hosting the experiment and for technical support to operate the magnet and cryogenics. We thank V. Burwitz, G. Hartner of the MPE PANTER X-ray test facility for providing the opportunity to calibrate the X-ray telescope and for assistance in collecting and analysing the characterization data. We acknowledge support from NSERC (Canada); MSE (Croatia); CEA (France); BMBF (Germany) under the grant numbers 05 CC2EEA/9 and 05 CC1RD1/0, DFG (Germany) under grant numbers HO 1400/7-1 and EXC-153, GSRT (Greece), NSRF: Heracleitus II; Agence Nationale de la Recherche (France) ANR-19-CE31-133 0024; the Spanish Agencia Nacional de Investigación (AEI) under grants PID2019-108122GB-C31 and PID2022-137268NB-C51,funded by MCIN/AEI/10.13039/501100011033/FEDER, as well as the “Planes complementarios, Programa de Astrofísica y Física de Altas Energías” from the “European Union NextGenerationEU/PRTR” funds; the European Research Council (ERC) under grants ERC-2017-AdG 788781 (IAXO+) and ERC 802836; the Marie Skłodowska-Curie grant agreement No 101026819 (LOBRES); Turkish Atomic Energy Authority (TAEK); NSF (USA) under Award No. 0239812; NASA under the grant number NAG5-10842; and IBS (Korea) with code IBS-R017-D1. Part of this work was performed under the auspices of the US Department of Energy by Lawrence Livermore National Laboratory under Contract No. DE-AC52-07NA27344 and funded through Lawrence Livermore National Laboratory Directed Research and Development project LDRD-17-ERD-030.